I need help with this equation

24x+3x2=0

Answer by checkley71(8403)

24X+3X^2=0
X(24+3X)=0
3X=-24
X=-24/3
X=-8 ANSWER.
PROOF
24*-8+3*-8^2=0
-192+3*64=0
-192+192=0
0=0